OpenStudy (anonymous):

Solve for x showing all of your work: −9x + 1 = −80

OpenStudy (anonymous):

Subtract 1 from both sides −9x=−80−1 Simplify −80−1 to −81 −9x=−81 Divide both sides by −9 x=−81/−9 Simplify 81/−9 to −81/9 x=81/9 Simplify 81/9 to 9 x=9

OpenStudy (anonymous):

Thanks again!

OpenStudy (anonymous):

yup